OpenStudy (cookiemonster18):

For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (graph) X 32 16 8 Y 24 8 6 yes; y=0.5x yes; y=2x yes; y=1.5x no; y does not vary directly with x.

OpenStudy (mathmale):

It's unusual for x-values to be listed in decreasing (instead of increasing) order. You may find the problem easier to understand if you re-write the table to show x in increasing order. x 8 16 32 y ?

OpenStudy (cookiemonster18):

y would be 6,8 ,24

OpenStudy (cookiemonster18):

@mathmale

OpenStudy (welshfella):

direct variation is described by y = kx where k is a constant value so k = y/x check if this is true for the values in your table.

OpenStudy (cookiemonster18):

I think the answer is C yes; y=1.5x am I correct? @welshfella

OpenStudy (welshfella):

work out the different values of y/x for example 6/8 = 0.75 8/16 = ?

OpenStudy (welshfella):

if they are different ten the answer is D

OpenStudy (cookiemonster18):

8/16=0.5 24/32=0.75 but how is it D?? @welshfella because all the answers to each of the fractions give different answers?

OpenStudy (welshfella):

because the value of k is not constant- so y does not vary directly with x

OpenStudy (welshfella):

yes